### v1

# Large Scale Multi-Task Bayesian Optimization with Large Language Models

Abstract
In multi-task Bayesian optimization, the goal is to leverage experience from optimizing existing tasks to improve the efficiency of optimizing new ones. While approaches using multi-task Gaussian processes or deep kernel transfer exist, the performance improvement is marginal when scaling to more than a moderate number of tasks. We introduce a novel approach leveraging large language models (LLMs) to learn from, and improve upon, previous optimization trajectories, scaling to approximately 2000 distinct tasks. Specifically, we propose an iterative framework in which an LLM is fine-tuned using the high quality solutions produced by BayesOpt to generate improved initializations that accelerate convergence for future optimization tasks based on previous search trajectories. We evaluate our method on two distinct domains: database query optimization and antimicrobial peptide design. Results demonstrate that our approach creates a positive feedback loop, where the LLM’s generated initializations gradually improve, leading to better optimization performance. As this feedback loop continues, we find that the LLM is eventually able to generate solutions to new tasks in just a few shots that are better than the solutions produced by “from scratch” by Bayesian optimization while simultaneously requiring significantly fewer oracle calls.
Machine Learning, ICML
1 Introduction
Multi-task optimization seeks to use related, previously solved tasks to accelerate the optimization of new ones. Multi-task optimization appears naturally in a variety of domains where similar problems are encountered repeatedly, such as hyperparameter optimization, material science, database query optimization, and drug design. Formally, suppose we have tasks $\{1,2,\ldots,T\}$, each associated with its own objective function $f_{t}(\mathbf{x})$. In multi-task Bayesian optimization (BO), we seek to find some $\mathbf{x}^{*}_{t}$ such that
$\mathbf{x}_{t}^{*}=\underset{\mathbf{x}\in\mathcal{X}}{\arg\min}f_{t}(\mathbf{x}),$ (1)
for each task $t$. We focus on the setting where, for each task, we have collected a dataset $D_{t}$ of observations, and we wish to leverage this data when optimizing unseen test tasks.
Multi-task BO methods have predominantly used Gaussian processes (GPs) and/or a variety of shared weight neural network feature extractors to jointly learn correlations across tasks (Swersky et al., 2013; Perrone et al., 2018; Patacchiola et al., 2020; Hakhamaneshi et al., 2022). A standard approach involves placing a multi-output GP over the input-task space, decomposing the kernel as an input kernel $k(\mathbf{x},\mathbf{x}^{\prime})$ and a task kernel $k(t,t^{\prime})$. Despite their effectiveness, many of these methods—with the notable exception of recent work such as Wang et al. (2024c)—tend to saturate in performance after tens of training tasks and do not extract additional optimization performance on new tasks when given hundreds or thousands of related tasks for training.
We propose Bayesian Optimization with LLM Transfer (BOLT), a straightforward approach to multi-task BO that departs from the framework of building related task information into the BO surrogate model. Instead, as BO completes optimization for training tasks, we fine-tune a large language model (LLM) to, given a task description or context $C[f_{t}]$, generate solutions for that optimization problem that we can use as strong initialization for BO.
This approach creates a self-reinforcing feedback loop: BO generates high-quality solutions that we can leverage to fine-tune the LLM; the fine-tuned LLM, in turn, produces better initializations that improve BO performance. Over time, the LLM learns to directly generate solutions that are highly competitive, enabling top-of-$k$-samples from the LLM (requiring just a few oracle calls) to outperform full “from scratch” BO runs (requiring a large number of oracle calls). This iterative improvement enables BOLT to scale to and still extract value from thousands of tasks.
We validate BOLT on two diverse and challenging domains where many related optimization tasks are available.
Application 1: Database query plan optimization. In database query plan optimization, the goal is to optimize execution plans in PostgreSQL to minimize runtime. Efficient query planning is critical for database performance, and traditional optimizers such as PostgreSQL’s built-in query planner rely on hand-crafted heuristics and cost models. We show that an LLM fine-tuned using BOLT can generate superior plans, sometimes outperforming PostgreSQL’s own optimizer in a single-shot setting.
Application 2: Antimicrobial peptide (AMP) design. Antimicrobial peptides are small peptides that kill (pathogenic) bacteria. We use a software package, APEX (Wan et al., 2024), which can predict the minimum inhibitory concentration (MIC) of peptides against various pathogens.
In this setting, each task is a distinct extinct peptide used as a template, and the goal is to enhance its antimicrobial activity through a limited number edits. We demonstrate that our approach continues to improve as more tasks are introduced and, in the few-shot setting, can eventually generate peptides that outperform those found through full “from-scratch” BO runs.
Our experimental results highlight two key benefits of BOLT. First, unlike many existing multi-task BO methods, LLM-generated initialization continues to improve performance as the number of tasks grows, avoiding the saturation observed in GP-based methods. Impressively, by the end of training, the LLM-generated initializations alone nearly outperform competing approaches. Second, a sufficiently fine-tuned LLM can eventually produce solutions in a few-shot setting that rival or surpass full “from scratch” BO runs. Notably, in the database query optimization setting, the LLM becomes a better one-shot optimizer than PostgreSQL’s built-in query planner for certain query types.
Contributions
  1. 1.
    We propose BOLT, a scalable and simple alternative to traditional multi-task BO, leveraging LLMs to generate strong initial solutions for new tasks. BOLT leverages a combination of high quality optimized solutions produced by BO and self augmentation for fine-tuning.
  2. 2.
    We validate BOLT through experiments on two challenging domains: database query optimization and antimicrobial peptide design. Results demonstrate that initializations generated by BOLT continue to improve performance as the number of tasks grows.
  3. 3.
    We provide empirical results demonstrating that, after sufficient fine-tuning, the LLM alone can eventually match or even outperform full “from scratch” Bayesian optimization runs with significantly fewer oracle calls.
2 Background
Bayesian optimization (BO).
Bayesian Optimization (BO) (Močkus, 1975; Snoek et al., 2012) is an iterative approach to optimize black-box functions in a sample-efficient manner. On each step of the optimization, a supervised probabilistic surrogate model (usually a Gaussian Process (GP) (Rasmussen, 2003)) is conditioned on all data collected so far. Then, the surrogate model’s predictive posterior distribution $p(y\mid\mathbf{x},D)$ is used to decide what data point(s) should be evaluated next, typically by maximizing some acquisition function, defined with respect to $p(y\mid\mathbf{x},D)$, which guides the exploration-exploitation trade off. Finally, the selected data points are evaluated on the black-box function and added to the dataset. This iterative process continues until the evaluation budget is reached.
Structured optimization via latent space BO. BO has recently been applied to optimizing structured search spaces, such as molecular and amino acid sequences, by leveraging latent space Bayesian optimization. This approach incorporates a variational autoencoder (VAE) to map structured inputs into a continuous latent space, where BO is performed (Eissman et al., 2018; Tripp et al., 2020; Grosnit et al., 2021; Siivola et al., 2021; Stanton et al., 2022; Maus et al., 2022). Structured inputs $\mathbf{x}$ (e.g., amino acid sequences) are mapped to continuous latent representations $\mathbf{z}$ by the VAE encoder $\Phi(\mathbf{x})$. This creates a transformed continuous (latent) representation of the structured search space where BO can be directly applied. The corresponding latent candidate points are then decoded by the VAE decoder, $\Gamma(\mathbf{z})$, to reconstruct structured outputs for evaluation. For large combinatorial structured search spaces, such as the space of organic molecules or the space of all peptide amino acid sequences, the latent space of the VAE is typically high-dimensional (on the order of several hundred dimensions) in order to represent the large structured space effectively.
Optimizing antimicrobial peptides. In antimicrobial peptide design, we seek peptides (sequences of amino acids) that minimize the MIC (minimum inhibitory concentration, measured in $\mu$ mol $\text{L}^{-1}$) for some target bacterial pathogen. MIC is a measure of the concentration of the peptide required to inhibit growth of the target bacterial pathogen (Kowalska-Krochmal & Dudek-Wicher, 2021). A key challenge in antimicrobial peptide design is that many modern bacterial pathogens have developed resistance to modern antibiotics. To solve this challenge, Wan et al. (2024) propose designing new peptides with high sequence similarity to template peptides mined from extinct organisms. The template peptides themselves do not typically achieve sufficiently low MIC for target bacteria pathogens. However, since these template peptides have not been encountered in nature for thousands for years, modern antimicrobial resistant bacteria have not evolved resistance to them. It follows that new peptides are more likely to evade antibiotic resistance if they are designed to be similar to the extinct template sequences. We employ this strategy, optimizing antimicrobial peptides with a minimum threshold sequence similarity to the extinct template peptides from Wan et al. (2024). We also employ latent space BO to optimize over the structured, combinatorial space of amino acid sequences.
Optimizing database query plans. Query optimization in data management systems involves translating a declarative SQL query into an execution plan that efficiently retrieves the correct results (Graefe & McKenna, 1993). This problem has been extensively researched in the field of data management (Selinger et al., 1989), as the difference in execution time between an optimal and a poorly chosen query plan can be several orders of magnitude (Leis et al., 2015). Since individual query plans are composed of discrete characteristics (e.g. join order trees), the search space of possible query plans is structured and combinatorial. We therefore employ latent space BO. We use the string representation for query plans proposed by Tao et al. (2025) to pre-train a VAE model that maps the structured space of query plans to a continuous latent space where BO can be applied.
Database query plan optimization with right-censored observations. In database query optimization, our black-box objective function measures the execution latency of the query plan. The latency difference between a “good” query plan and a “bad” one can be orders of magnitude (Leis et al., 2015). This can lead to the majority of optimization runtime being taken up by evaluating a small number of poorly performing plans. A natural solution to this problem is to timeout objective function evaluations after they have reached some threshold latency $\tau$, resulting in right-censored observations. A right-censored observation is an observation at data point $\mathbf{x}$ where we observe only that $y\geq\tau$ for some chosen timeout threshold $\tau$, rather than observing the typical noisy objective value $y$. Prior work has been done to extend Bayesian optimization methods to the setting of right-censored observations. Hutter et al. (2013); Eggensperger et al. (2020) extended Bayesian optimization methods to the setting of right-censored observations by introducing an EM-like algorithm to impute the values of censored observations. Eggensperger et al. (2020) expanded on this, defining a single surrogate model capable of being conditioned on the combination of censored and uncensored data gathered during optimization.
Tao et al. (2025) extend this to the setting of approximate GP surrogate models. Since we focus on tasks that involve large function evaluation budgets, we employ Tao et al. (2025)’s proposed method of modeling censored data with approximate GPs.
3 Bayesian Optimization with LLM Transfer (BOLT)
We propose an iterative framework for using large language models (LLMs) to improve Bayesian optimization (BO) performance across a family of related tasks. We are given a set of $T$ training tasks defined by objective functions $f_{1}(\mathbf{x}),...,f_{T}(\mathbf{x})$. We additionally assume that, for each objective function we have a context or task description $C[f_{t}]$ that can be a natural language or other input description that differentiates $f_{t}$ from any other task in the application domain. For example, this might be a SQL query we are trying to optimize.
For each training task, we assume we have optimized the objective with some BO procedure, resulting in the optimization trajectories $\{\mathcal{D}^{\star}_{t}\}_{t=1}^{T}$, with each $\mathcal{D}^{\star}_{t}$ containing the top-$K$ observations from the trajectory for the $t^{\text{th}}$ task.
Our goal is to leverage this training data to learn an LLM-based “initialization policy” $\pi$ that, when presented with new related tasks $\left\{f_{T+1}(\mathbf{x}),C[f_{T+1}(\mathbf{x})]\right\}$, proposes a high-quality set of candidate solutions for BO to further refine.
These two procedures—(1) using BO to collect high-quality data for training tasks, and (2) using the LLM to initialize BO for new tasks—can be used as an “outer-loop”/“inner-loop” approach to solving a large number of related tasks sequentially, where the LLM is periodically updated as more optimization runs complete.
Because the LLM and BO only interact through generating initialization and generating fine tuning data respectively, our approach here is relatively agnostic to the specific underlying implementation of BO used to optimize for each task. This enables the straightforward use of the full range of recent BO advances on multi-task, constrained, and other optimization settings.
3.1 Initializing BOLT.
At initialization for a workload of tasks, we have only an un-tuned LLM BOLT-00 that is generally useless for the task setting because it is unaware of even the specific format for candidate suggestions. For the first iteration, we solve $T$ optimization tasks with a single-task Bayesian optimization routine where we initialize BO using some standard initialization procedure. We run optimization on each of the $T$ initial tasks, and extract the optimization trajectories $\{\mathcal{D}^{*}_{i}\}_{i=1}^{T}$ from each run.
The number of tasks $T$ that we collect at initialization time–as well as during each round of the BOLT “outer-loop” represents a non-trivial trade-off due to the computational cost of both running BO and the cost of fine-tuning the LLM. Fine-tuning the LLM more frequently results in both additional computational and monetary costs, but allows subsequent BO runs to complete more efficiently.
In this paper, we generally erred on the side of lower monetary cost in exchange for additional compute cost. Specifically, while processing an entire workload of training and test tasks, we only fine-tune the LLM three times.
3.2 LLM Fine-Tuning
The fine-tuning process employs supervised learning using OpenAI’s GPT-4-mini-0718 model through their API. From the optimization trajectories $\{\mathcal{D}^{*}_{i}\}_{i=1}^{T}$, we extract the top-$K$ observations from each of the $T$ runs completed so far. We use these observations along with the task contexts $\left\{C[f_{t}]\right\}_{t=1}^{T}$ to construct a fine-tuning data $\mathcal{D_{\textrm{ft}}}$. Each training instance contains three components:
  1. 1.
    A system prompt shared across all tasks in the workload/problem domain, which specifies the objective (e.g., generating efficient join orderings).
  2. 2.
    A user prompt containing the task-specific context $C[f_{t}]$ (e.g., the SQL query requiring optimization).
  3. 3.
    A response prompt containing the high-performing solution $x$ discovered through BO.
{minted} [breaklines, fontsize=]latex System: You are a helpful assistant that provides efficient join orderings for given queries. User: SQL query to be optimized Assistant: Optimized query plan
Figure 1: The prompt template used for prompting gpt-4o-mini for generating optimized query plans.
We fine-tune using OpenAI’s standard fine-tuning API (OpenAI et al., 2024). Specifically, we format our data into the required JSONL format (i.e., prompt-solution pairs) and then upload it via the fine-tuning API to initiate training. The model is trained to minimize the negative log-likelihood of the solution tokens x given the task context $C$:
$\mathcal{L}=-\sum_{i=1}^{|\textbf{x}|}\log\pi(x_{i}|C,\textbf{x}_{<i})$ (2)
We note that our approach leverages full model fine-tuning rather than extensive prompt engineering (Lester et al., 2021; Li & Liang, 2021). This allows the model to learn the task requirements through the context-solution pairs in $\mathcal{D_{\textrm{ft}}}$, rather than explicit instructions. However, for scenarios requiring few-shot learning on untrained models, more careful prompt engineering may be beneficial.
This fine-tuning process produces an updated model that encodes the knowledge from $\mathcal{D_{\textrm{ft}}}$. In our experiments, we will refer to an LLM trained on $T$ tasks in this way as BOLT-$T$.
3.3 Using the LLM for Multi-Task BO
Once we have a fine‐tuned model, BOLT-$T$, we can leverage the fine-tuned LLM’s capabilities to generate higher-quality initialization points for subsequent optimization tasks.
For a set of $n$ new tasks $\{t_{i}\}_{i=T+1}^{T+n}$, we sample from BOLT-$T$ to generate the same number of initialization points used by the baseline “from scratch” approach. The sampling prompt maintains the same structure as the training prompt without the assistant response. The BOLT-$T$ generated solutions are refined with a standard BO routine, and the top‐$K$ performing solutions for each task $t$—along with their contexts $C[f_{t}]$—are incorporated into the training set for the next round of fine‐tuning.
Closed-loop feedback. This iterative process creates a positive feedback loop between BO and the LLM. After completing each iteration of BO, we augment the LLM fine-tuning dataset with the newly collected high-scoring solutions and their corresponding contexts. Re-training the LLM on this expanded dataset allows it to internalize additional insights over the tasks and produce more effective initialization for the next batch of tasks. Through multiple iterations, the LLM learns to propose increasingly high-quality starting points, thereby jump-starting subsequent BO runs and reducing the time needed to find near‐optimal solutions. Ultimately, this positive feedback loop continually refines the model’s understanding of what a good solution should look like under each context, improving its performance across the entire family of tasks.
Algorithm 1 Inner Loop: LLM-Initialized Bayesian Optimization
0:  Task $t$, context $C[f_{t}]$, fine-tuned LLM $\pi_{n}$, evaluation budget $B$, batch size $b$
0:  Optimized solutions $X^{*}_{t}$
1:  $X_{\text{init}}\leftarrow\pi_{n}(C[f_{t}])$LLM proposes initial candidates}
2:  Evaluate $y_{\text{init}}\leftarrow f_{t}(X_{\text{init}})$
3:  Initial dataset $\mathcal{D}\leftarrow\left(X_{\text{init}},y_{\text{init}}\right)$
4:  Initialize GP surrogate model $\mathcal{GP}(X_{\text{init}},y_{\text{init}})$
5:  for step $i=1$ to $\lfloor B/b\rfloor$ do
6:     Select $X_{\text{next}}\leftarrow\operatorname*{arg\,max}_{\alpha}(x;\mathcal{GP})$Acquire new candidates}
7:     Evaluate $y_{\text{next}}\leftarrow f_{t}(X_{\text{next}})$
8:     $\mathcal{D}\leftarrow\left(X\cup X_{next},y\cup y_{next}\right)$Update dataset}
9:     Update $\mathcal{GP}$ with new observations
10:  end for
11:  Return $X^{*}_{t}\leftarrow\text{top-}K(X)$Best solutions for LLM fine-tuning}
Algorithm 2 Outer Loop: LLM Fine-Tuning via BO Trajectories
0:  Initial dataset $\mathcal{D}_{0}=\{(C[f_{t}],x_{i},y_{i})\}$, initial LLM $\pi_{0}$, iteration count $T$
0:  Fine-tuned LLM $\pi_{T}$
1:  Initialize $\mathcal{D}\leftarrow\mathcal{D}_{0}$, $\pi\leftarrow\pi_{0}$
2:  for iteration $k=1$ to $T$ do
3:     for each task $t$ in batch do
4:        $X^{*}_{t_{j}}\leftarrow$ InnerLoop($t$, $\pi_{k}$, $B$, $b$) {BO with LLM initialization}
5:        $\mathcal{D}\leftarrow\mathcal{D}\cup\{(C[f_{t}],x,y)\,|\,x\in X^{*}_{t}\}$Augment with top solutions}
6:     end for
7:     Fine-tune $\pi_{k}$ on $\mathcal{D}$ using instruction prompting
8:  end for
3.4 Self-Augmentation
As the fine-tuned LLM enhances few-shot generation with more optimization data, it is worth exploring whether the costly sequential BO processes can be minimized. Thus, we explore “self-improvement” methods to refine the LLM policy without the expense of additional optimization runs (Algorithm 3).
Specifically, once an LLM has been fine-tuned using some of the tasks we set aside for training, we prompt it to generate additional solutions for all available tasks in that problem setting. We then score these solutions using the problem’s oracle and fine-tune the LLM again directly with this labeled self-generated data, in a manner similar to self-play in reinforcement learning or self-instruction in LLM training (Haluptzok et al., 2022; Shypula et al., 2024). By filtering and fine-tuning on its own best outputs, the LLM can iteratively teach itself how to propose better solutions.
Algorithm 3 Self-augmentation for LLM Finetuning
0:  Set of tasks $\mathcal{T}$, fine-tuned LLM $\pi_{\theta}$, iteration count $T$, criteria $\mathcal{C}$
0:  Fine-tuned LLM $\pi_{\theta+T}$
1:  for iteration $k=1$ to $T$ do
2:     for task $t\in\mathcal{T}$ do
3:        $X_{init}\sim\pi_{\theta+k}(t)$Generate samples from optimized LLM $\pi_{\theta+k}$}
4:        $X^{\star}_{init}\leftarrow\mathcal{C}\{x\,|\,x\in\mathcal{X}^{\star}_{init}\}$Select best samples using the criteria for further fine-tuning}
5:     end for
6:     $\mathcal{D}\leftarrow\mathcal{D}\cup\{(C[f_{t}],x,y)\,|\,x\in X^{*}_{init}\}$Augment with top solutions}
7:     Fine-tune $\pi_{\theta+k}$ on $\mathcal{D}$ using instruction prompting
8:  end for
4 Experiments
We evaluate BOLT on two distinct problem domains, each with a large number of similar tasks. Both domains allow their problem definitions and solutions to be represented as strings. This allows BOLT to operate both in sequence space, where the LLM learns from optimization trajectories, and in latent space, where Bayesian optimization (BO) makes additional progress using LLM-sampled initializations.
4.1 Experimental Setup
Implementation details.
For the inner optimization loop, we implement a constrained version of the LOL-BO algorithm (Maus et al., 2022) using BoTorch and GPyTorch (Balandat et al., 2020; Gardner et al., 2018). Code to reproduce all results will be publicly available on GitHub after acceptance, with a link to the repository at placeholder. For query optimization, we use an acquisition batch size of 1 with a budget of 4,000 oracle calls, while for peptide design, we employ a larger acquisition batch size of 50 with a budget of 200,000 oracle calls.
The outer loop uses instruction prompting (Mishra et al., 2021; Longpre et al., 2023) to guide the LLM in producing optimized sequences. Figure 1 shows the prompt used to prompt gpt-4o-mini for efficient query plans. After each optimization iteration, we augment the training set with the highest-scoring sequences from the optimization trajectory and fine-tune gpt-4o-mini on this expanded dataset. When fine-tuning the LLM for the query plan optimization task, we use OpenAI’s automatic batch size selection option. For the peptide design task, we found that using the automatic batch size option did not provide a similar boost in performance, and we use a constant batch size of $10$. For both tasks, we fine-tuned the LLM for $2$ epochs and used the default OpenAI LR multiplier hyperparameter of $1.8$. To ensure the solutions always have the correct syntax, we filter out characters that do not correspond to strings of integers or valid amino acids for the respective tasks. Appendix A provides additional details on the fine-tuning process and prompts used for the peptide task.
Query optimization task. This task focuses on finding query plans (including join orderings and their operators) that minimize execution time for a given query. We take a subset of $2933$ queries from the Cardinality Estimation Benchmark introduced by Negi et al. (2021), keeping $99$ queries for validation. Following Tao et al. (2025), we allow BO over query plans by encoding join orders and operators as integer lists, which are then mapped to a 64-dimensional continuous latent space using the pre-trained query plan VAE from Tao et al. (2025). For the pretraining dataset, we randomly generate $1,169,890$ query plans generated based on the database schema, separated into 80/10/10 splits. For the initial “from scratch” runs with no LLM, we initialize with the set of 50 query plans used by BAO (Marcus et al., 2021), a default query optimizer that produces reasonable but non-optimal plans. Subsequent runs use 50 LLM-sampled query plans per query as initialization points. The “task description context” $C$ used to fine-tune the LLM for this task is the full SQL query string.
Antimicrobial peptide design task. For this application, we are given a library of 1000 extinct, weakly antimicrobial seed peptides $S=\left\{s_{1},...,s_{L}\right\}$. A task in this setting is to take a particular seed peptide $s_{i}$ and make modifications to it to minimize the minimum inhibitory concentration (MIC) against A. Baumannii ATCC $19606$, measured in $\mu\text{mol}/L$. We created a library of $L=1000$ extinct peptides and held out the last $100$ as validation. We ensure edited peptides maintain a minimum $75\%$ similarity to the seed peptide, defined by $1-\frac{d(S,S^{\prime})}{\text{len}(S)}$, where $d$ is the Levenshtein distance between them. All of the validation peptides are at least $25\%$ different from any other peptide in the library. Although the seed peptides don’t achieve low MICs, the hope is that bacteria are less likely to have developed resistance to their variations as they come from extinct species (Wan et al., 2024). We assess MICs with the APEX model Wan et al. (2024) and utilize a VAE trained on 4.5 million amino acid sequences Torres et al. (2024) to map peptides into a $256$-dimensional latent space. Initial optimization uses $1000$ randomly mutated sequences with a similarity constraint of $75\%$ to the seed. Subsequent runs utilize $1000$ LLM sampled peptides. We use the seed amino acid sequence as the “task description context” $C$ for LLM fine-tuning.
Baselines. We compare BOLT against two baseline approaches. First, we compare to “from scratch,” single task LOL-BO which we will refer to as STBO, which operates without prior task knowledge. Second, we compare to a common strategy for multi task BO—see, e.g., Patacchiola et al. (2020); Hakhamaneshi et al. (2022); Perrone et al. (2018), where a shared GP is trained on all tasks through a neural network feature extractor using the optimization trajectories from training tasks. This shared GP is then check-pointed and used on test tasks. Several papers have found success with variations of this approach. ABLR (Perrone et al., 2018) uses independent Bayesian linear regression heads per task on the shared feature extractor, while FSBO (Wistuba & Grabocka, 2021) uses an adaptation of DKT (Patacchiola et al., 2020). In this paper, for the final supervised model, we use the same PPGPR model as the BO inner-loop in our method as we require scalability but find this results in better performance than Bayesian linear regression or random Fourier feature models (Rahimi & Recht, 2007). Additional details can be found in Section B.1.
4.2 Optimization Results
Initializing Bayesian optimization with BOLT. To initialize BO with BOLT, we sample 50 points for query optimization tasks and 1,000 points for peptide optimization tasks. All sampling is performed with a temperature parameter of 0.7 for the query optimization task and 1.0 for the peptide design task unless otherwise specified. In Figure 2, we demonstrate that initializing BO with BOLT significantly improves optimization efficiency across both domains. On the query optimization task (left), while DKT/FSBO makes improvements over STBO, the gains appear to plateau after only 20 tasks. In contrast, BOLT successfully scales to over 1400 tasks and converges to far higher quality solutions faster. On the peptide design task (right), BOLT shows similarly strong performance, while DKT/FSBO struggles to take advantage of the data collected for separate templates. Surprisingly, we find that on both tasks once BOLT reaches a sufficient scale, it begins to generate initialization data for BO that is significantly better performing than the final results found by STBO and DKT/FSBO.
Figure 2: Bayesian optimization performance on (Left) query plan optimization and (Right) antimicrobial peptide design. In both settings, BOLT outperforms or matches baselines with just initialization data before optimization begins.
BOLT as a one- and few-shot optimizer. As illustrated in Figure 3, BOLT demonstrates strong few-shot generalization capabilities, even achieving single-shot performance competitive with traditional systems.
In query optimization, all BOLT variants outperform the top BAO solution within 5 samples. Notably, BOLT-1138 and BOLT-1426 surpass PostgreSQL in a single sample, indicating their potential for rapid deployment in low latency scenarios. The performance of BOLT consistently improves with more iterations across both tasks, except at 50 samples on the query optimization task, where BOLT‐1138 slightly outperforms BOLT‐1426. This may be due to variances in LLM sample generation or training. Overall, these findings confirm our approach’s robustness and efficiency when scaling BOLT to thousands of tasks.
We further compare our few‐shot performance (Figure 3, before the x-axis break) to full BO runs (Figure 3 after the x-axis break). In both tasks, BOLT achieves few shot results comparable to the full BO runs.
Figure 3: Evaluating BOLT in the few shot setting and comparing to full optimization runs in both problem settings (Left: query plan optimization; Right: peptide design). In each plot, we show objective values accumulated across all validation tasks for various methods. Before the break in each plot, we show the few shot performance of BOLT and relevant domain specific few shot baselines (the PostgreSQL one-shot query optimizer, Bao). After the break in each plot, we show the results at the end of various full BO runs, which are often comparable to BOLT’s few shot performance.
4.3 Ablation Studies
In this section, we provide additional experiments analyzing various components and extensions of BOLT.
Improving LLM performance via self-augmentation. We investigate whether the performance improvements of self-augmentation as outlined in Algorithm 3 can improve the LLM while avoiding the computational expense of inner-loop BO. We apply the self-augmentation process to BOLT-1138 and BOLT-1426, generating 10 samples from each across all $2,933$ training tasks, keeping only queries that outperform the best query plan from BAO’s solutions. We then use these two dataset as additional fine-tuning to create self-augmented versions of the models.
Table 2 shows that self-augmentation yields substantial improvements even without additional tasks optimized by BO. Both self-augmented models converge to a similar performance level, achieving a summed runtime of about 62 seconds across the 100 validation queries. This convergence suggests a natural performance plateau after training on either 1,500 tasks (BOLT-1426) or 1,100 tasks plus self-generated samples (BOLT-1138+SA). The consistency of this plateau across different training approaches further demonstrates BOLT’s robustness when scaling to large task sets. This self-augmentation experiment indicates that once the LLM has been fine-tuned to sufficient performance, it can generate additional fine-tuning data, reducing the number of actual BO optimization runs required. Additionally, our framework scales well with more training tasks without performance loss.
Ablating the impact of temperature on performance. In all prior experiments, we used a constant sampling temperature of 0.7 for response generation. We demonstrate that greedy decoding (i.e., setting the temperature to 0) improves one-shot performance. As shown in Figure 4, this strategy generates a larger set of queries that surpasses the best among 50 BAO-generated queries by a wider margin.
Figure 4: Comparing one-shot LLM generated initialization to the best of the 50 BAO plans. (Top): One-shot generation from BOLT-$1426$, sampling temperature $T=0.7$. (Bottom): One-shot generation from BOLT-$1426$, greedy sampling $T=0.0$. In the one-shot setting, purely greedy sampling is better (lower, more blue).
Best@ BOLT-1138 BOLT-1138 BOLT-1426
Best@50 78.16 64.03 63.68
Best@20 82.59 70.52 66.23
Best@10 90.19 74.40 70.21
Best@5 102.99 85.21 76.28
Best@2 127.97 129.26 102.29
Best@1 202.04 193.64 160.22
Table 1: Comparing LLMs fine tuned with (Left) data from 1138 tasks, (Right) data from 1426 tasks, and (Middle) data from 1138 tasks, but including the extra data from BOLT-1426, and removing data from older tasks. See ablation study #3 for details.
Ablating the impact of data quality on training. We perform an ablation to assess the importance of using ”better” versus ”more” training data for fine-tuning LLMs through iterations of BOLT. Starting with the BOLT-$1138$ model, we collect top solutions from a new BO round and train two variants: 1) BOLT-$1426$, which adds all new solutions to the original BOLT-$1138$ set. 2) BOLT-$1138^{\star}$, which instead replaces an equal number of old solutions to maintain the same training set size. As shown in Table 1, both benefit from higher-quality data, suggesting “better” data significantly boosts performance. However, BOLT-$1138^{\star}$ underperforms BOLT-$1426$, which incorporates more and better data, confirming that both factors enhance model performance.
Method Best@50
LLM (BOLT-50) 87.84
LLM (BOLT-893) 82.31
LLM (BOLT-1138) 78.16
LLM (BOLT-1426) 63.68
LLM (BOLT-50) + SA 82.25
LLM (BOLT-893) + SA 63.05
LLM (BOLT-1138) + SA 61.46
LLM (BOLT-1426) + SA 61.54
Table 2: Ablation study for self augmentation (SA) conducted on the query optimization task. For each of two LLMs with different training task sizes, we perform SA and generate 50 query plans from the LLM. We measure the best summed query execution time across the validation tasks from among these 50 samples.
5 Related Work
Language models as optimizers. Large language models (LLMs) have recently gained attention as sequence optimizers capable of tackling diverse black-box tasks where direct gradient information is unavailable or difficult to compute. LLM-based optimizers leverage the flexibility of natural language prompts to encode candidate solutions, constraints, and relevant task information. Methods like OPRO illustrates how iterative prompting can refine solutions (Yang et al., 2024; Zelikman et al., 2024), while other approaches integrate self-improving strategies that reuse high-performing LLM outputs for further fine-tuning (Shypula et al., 2024). This set of techniques has been applied to biophysical domains such as molecular design and protein engineering, where the LLM proposes mutations to enhance certain properties, as well as to program optimization tasks where the LLM speeds up code execution time (Shypula et al., 2024; Wang et al., 2024a; Madani et al., 2023).
Database optimization. Recent work has applied Bayesian optimization (BO) to improve overall database performance (Zhang et al., 2022; Nardi et al., 2019; Cereda et al., 2021) by tuning the parameters of the database configuration. As far as we are aware, Tao et al. (2025) were the first to apply BO to the specific setting of database query plan optimization considered in this paper. Other work has applied Reinforcement Learning (RL) to database query plan optimization (Marcus et al., 2019; Yang et al., 2022; Zhu et al., 2023). RL query optimizers learn from mistakes and improve performance over time. Unlike BO, however, RL requires large supervised datasets for pre-training and typically aims to minimize cumulative query latency rather than achieving the lowest possible latency.
6 Discussion and Limitations
We first highlight a few limitations. First, our approach not only requires that all tasks in a problem setting have the same input domain (a problem that has been explored e.g. by Fan et al. (2022)). We further require the existence of a task description context $C[f_{t}]$ that can be used in an LLM prompt to define the task. This is likely more difficult, e.g., for hyperparameter optimization, where the primary thing distinguishing tasks is the data that the models are to be trained on; for this setting, approaches such as (Wang et al., 2024b) are likely more appropriate. Finally, we note that the cost of LLM fine-tuning is significantly higher than simple gradient updates of a shared feature extractor.
Despite these limitations, in two real-world applications where BOLT was applicable it yielded strong results. Few-shot generation matched “from scratch” BO runs, and initializing BO from the LLM samples often improved performance further. Moreover, the interplay between the LLM and Bayesian optimization is noteworthy. Despite interest in using LLMs for optimization (Yang et al., 2024; Zelikman et al., 2024; Shypula et al., 2024; Wang et al., 2024a; Madani et al., 2023), finding initial strong solutions to fine-tune them is challenging in some domains. Bayesian optimization, by offering in-depth search, is an excellent candidate for this.
Impact Statement
This work includes an application in designing peptides using AI, specifically for combating antimicrobial-resistant bacteria—a critical challenge in modern medicine. While our work aims to accelerate the development of peptides that can address the growing threat of antimicrobial resistance, we acknowledge the potential for misuse of any “AI for drug design” method to develop harmful biological agents rather than useful drugs (Urbina et al., 2020).
Appendix A Prompt for antimicrobial peptide design.
{minted} [breaklines, fontsize=]latex System: You are a specialized assistant that modifies peptide sequences to enhance antimicrobial activity. Make up to 25User: Seed peptide to be modified Assistant: Modified peptide
Figure 5: The prompt template used for prompting gpt-4o-mini for generating optimized peptide sequences.
Appendix B Training details.
B.1 DKT/FSBO implementation details
For the antimicrobial peptide design task, we train a PPGPR model using the GPyTorch module with a fully connected network with two hidden layers, each with a dimension of 256. We use a batch size of 128, a learning rate of 0.01, and 1024 inducing points for all peptide design experiments.
For the database query plan optimization task, we instead use a fully connected network with two hidden layers, each with a dimension of 64. We use a batch size of 16, a learning rate of 0.01, and 1024 inducing points for all database query plan optimization experiments. Both models were trained for 20 epochs.
B.2 LLM self-augmentation details
For the antimicrobial peptide design task, in each round of self-augmentation, we generate 200 samples for each of the first $800$ train peptides, and add all peptides with a predicted MIC that is lower than $8$ (signaling significant antimicrobial activity) to the training set for the next round of BOLT.
For the database query plan optimization task, we generate 10 samples for each of the 2933 train queries and add all query plans with a running time that is lower than the best plan generated by BAO to the training set for the next round of BOLT.